\begin{thebibliography}{10}
\bibitem{H}
W. Heisenberg,
{\em Uber den anschaulichen Inhalt der quantentheoretischen Kinematik und Mechanik},  Zeitschrift for Physik, {\bf 43}, 172 -- 98 (1927).


\bibitem{M}
A. Messiah, {\em Quantum mechanics}, vol. I, North Holland Publ. Co., Amsterdam, 1962.


\bibitem{R}
H. P. Robertson, {\em The Uncertainty Principle}, Phys. Rev., {\bf 34}, 163, (1929).


\bibitem{S}
E. Schr\"{o}dinger, {\em Zum Heisenberschen Unsch\"{a}fprinzip}, Sitzungsber. Preuss. Akad. Wiss., {\bf XIX}, 296 -- 303, (1930)
[Proceedings of The Prussian Academy of Sciences.  Physics--Mathematical Section, {\bf XIX}, 296-303, (1930)]; English translation: {\em About Heisenberg Uncertainty Relation},
Bulg. J. Phys. {\bf  26}, 193--203, (1999); arXiv: 9903100 [quant--ph].

\bibitem{Mac}
L. Maccone, A. K. Pati,  {\em Stronger Uncertainty Relations for All Incompatible Observables}, Phys. Rev. Lett., \textbf{113}, 260401, (2014);\\
https://doi.org/10.1103/PhysRevLett.113.260401.


\bibitem{Hua}
Yichen Huang, {\em Variance-based uncertainty relations}, Physical Review A {\bf 86}, 024101, (2012); doi: 10.1103/PhysRevA.86.024101.


\bibitem{Bin}
Bin Chen,  Shao-Ming Fei, {\em Sum uncertainty relations
for arbitrary N incompatible
observables}, Scientific Reports {\bf  5}, 14238, (2015);  doi: 10.1038/srep14238.


\bibitem{Xiao}
Yunlong Xiao, Naihuan Jing, Xianqing Li-Jost and  Shao-Ming Fei,
{\em Weighted Uncertainty Relations}, Scientific  Reports {\bf 6}, 23201 (2016); doi: 10.1038/srep23201 (2016).


\bibitem{Pat}
A. K. Pati, P .K. Sahu, {\em Sum uncertainty relation in quantum theory}, Physics Letters A {\bf 367}, 177, (2007); doi:10.1016/j.physleta.2007.03.005.


\bibitem{Hof}
H. F. Hofmann and S. Takeuchi, {\em Violation of local uncertainty  relation as a signature of entaglement}, Phys. Rev. A {\bf 68}, 032103, (2003); doi: 10.1103/PhysRevA.68.032103.

\bibitem{Jam}
Farzad Dadipoura, Mohammad Sal Moslehiana, John M. Rassiasb, Sin--Ei Takahasi,
{\em Characterization of a generalized triangle inequality in normed spaces}, Nonlinear Analysis {\bf 75}, 735–--741, (2012); doi:10.1016/j.na.2011.09.004.


\bibitem{ku-rev1}
K. Urbanowski. {\em On a certain inequality for the sum of norms
and reverse uncertainty relations}, arXiv: 2503.05669v2 [quant-ph]

\bibitem{Bar}
S. M. Barnett and P. L. Knight, {\em Squeezing in correlated quantum systems}, Journal of Modern Optics,  {\bf 34}, 841 --- 853, (1987).
https://doi.org/10.1080/09500348714550781.


\bibitem{Cer}
P. Cerone and S. S. Dragomir, {\em  Mathematical Inequalities}
(Chapman and Hall/CRC, London, 2011).


\bibitem{Hsu}
Chih--Yung Hsu, Sen-Yen Shawb and  Han--Jin Wong, {\em Refinements of generalized triangle inequalities}, Journal of  Mathematical Analysis and Applications, {\bf  344},  17 (2008).



\bibitem{Son}
Qiu--Cheng Song, Jun--Li Li, Guang--Xiong Peng and  Cong-Feng Qiao,
{\em A Stronger Multi-observable Uncertainty Relation}, Scientific Reports {\bf 7}, 44764 (2017);
doi: 10.1038/srep44764.

\bibitem{strong-2}
Yao Yao, Xing Xiao, Xiaoguang Wang, and C. P. Sun,
{\em Implications and applications of the variance-based uncertainty equalities}, Physical Review A {\bf 91}, 062113, (2015); DOI: 10.1103/PhysRevA.91.062113.


\bibitem{Deb}
Debasis Mondal, Shrobona Bagchi, and Arun Kumar Pati, {\em Tighter uncertainty and reverse uncertainty relations},
Physical Review A {\bf 95}, 052117, (2017); DOI: 10.1103/PhysRevA.95.052117.


\bibitem{Xiao2}	
Yunlong Xiao, Naihuan Jing, Bing Yu, Shao-Ming Fei and Xianqing Li--Jost, {\em Near--Optimal Variance-Based
Uncertainty Relations}, Frontiers in Physics, {\bf 10},  Article No: 846330, (2022); DOI: 10.3389/fphy.2022.846330.


\bibitem{Xiao3}
Lei Xiao, Bowen Fan, Kunkun Wang, Arun Kumar Pati, and Peng Xue, {\em Direct experimental test of forward and reverse uncertainty relations },
Physical Review Research {\bf 2}, 023106 (2020); DOI: 10.1103/PhysRevResearch.2.023106.

\bibitem{ku-edu}
K. Urbanowski, {\em On the Weak Point of the Stronger Uncertainty Relation}, Academia Quantum, 2025, vol. 2, No. 1, (Academia.edu Journals);  doi:10.20935/AcadQuant7557.


\bibitem{ku-min}
K. Urbanowski, {\em On some states minimizing uncertainty
relations: A new look at these relations}, arXiv:2411.08131v4 [quant-ph].

\bibitem{ku1}
K. Urbanowski, {\em Uncertainty Relations: Curiosities
	and Inconsistencies}, Symmetry, {\bf 12}, 1640, (2020); doi: 10.3390/sym12101640.



\bibitem{Poz}
Victor Pozsgay, Flavien Hirsch, Cyril Branciard and Nicolas Brunner, {\em Covariance Bell inequalities}, PhysicaL Review A {\bf 96}, 062128 (2017): doi: 10.1103/PhysRevA.96.062128.


\bibitem{Khr}
Andrei Khrennikov and  Irina Basieva, {\em Entanglement of Observables: Quantum Conditional
Probability Approach}, Foundations of Physics, (2023) 53:84 (2023);
DOI: 10.1007/s10701-023-00725-7.




\bibitem{Xiao4}
Xiao Zheng, Ai--Ling Ji and Guo--Feng Zhang, {\em  Stronger reverse uncertainty relation for multiple incompatible observables},
Physica Scripta, {\bf 98}, 065113, (2023); DOI: 10.1088/1402-4896/acd484.





\bibitem{Dun}
C. F. Dunkl and K. S. Williams, {\em A Simple Norm Inequality},
The American Mathematical Monthly, {\bf 71}, No. 1,
53 --- 54, 1964.





\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{url}
\usepackage{amsfonts}
\usepackage[T1]{fontenc}
\usepackage{float}
\usepackage[utf8]{inputenc}
\usepackage{subfigure}
\usepackage{epsfig,epsf}
\usepackage{amsbsy}
\usepackage{amsfonts}
\usepackage{textcomp}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{graphicx}
\usepackage{amsmath,amssymb}
\usepackage{latexsym}
\usepackage{xcolor}
\usepackage{color}

\begin{document}

\title{Closer look at  sum  uncertainty relations\\ and related relations}
\author{
Krzysztof  Urbanowski\\
Institute of Physics,
University of Zielona G\'{o}ra,  \\
ul. Prof. Z. Szafrana 4a,
65-516 Zielona G\'{o}ra,
Poland\\
\hfill\\
e--mail:  K.Urbanowski@if.uz.zgora.pl, $\;$ k.a.urbanowski@gmail.com}

\maketitle

\begin{abstract}

We analyze the weak and critical points of various uncertainty relations that follow from the inequalities for the norms of vectors in the Hilbert space of states of a quantum system. There are studied uncertainty relations for sums of standard deviations, for sums of variances, and other relations between standard deviations or variances. The obtained results are compared with the conclusions obtained in similar cases using the standard Heisenberg--Robertson uncertainty relation. We also show that there  exists  an upper bound on the product of standard deviations that appears in the Heisenberg--Robertson uncertainty relation.

\end{abstract}
	
\section{Introduction }
In classical physics describing physical phenomena in our macroscopic world, physical quantities are described by functions in which the variables are positions and velocities (or momentum). The accuracy with which we measure these physical quantities depends only on the perfection of our measuring instruments and is independent of whether we simultaneously measure one or more physical quantities. In turn, in quantum mechanics describing processes occurring in the microscopic world, physical quantities are represented by observables, i.e. self--adjoint operators acting in the state space ${\cal H}$ of a given physical system.
And here, unlike the macroscopic world, if physical quantities under studies
are represented by non--commutative operators (say $A$ and $B$), $AB \neq BA$,
then unlike the classical world, it is not possible to unequivocally state what values the observables $A$ and $B$ will take when we try to simultaneously measure their values. This effect explains
the famous uncertainty principle discovered by Heisenberg. This principle belongs to the one of characteristic and the most important consequences of the quantum mechanics and indicates the fundamental differences between classical and quantum world.

Heisenberg's uncertainty principle states that the following relation must hold for the position and momentum \cite{H}:
\begin{equation}
\Delta_{\phi} x\ \cdot \Delta_{\phi} p_{x}\,\geq \,\frac{\hbar}{2}, \label{H1}
\end{equation}
where $\Delta_{\phi} x$ and $\Delta_{\phi} p_{x}$ are the standard (root--mean--square) deviations \cite{M}. The inequality (\ref{H1}) is the most known form of  the uncertainty principle and it was generalized by Robertson \cite{R} for two any non--commuting observables $A$ and $B$  and then by Schr\"{o}dinger \cite{S},
\begin{equation}
\Delta_{\phi} A \cdot \Delta_{\phi} B\;\geq\;\frac{1}{2} \left|\langle [A,B] \rangle_{\phi} \right|.\label{R1}
\end{equation}
Here
\begin{equation}
\Delta_{\phi} F \equiv \sqrt{\langle F^{2}\rangle_{\phi} - \langle F\rangle_{\phi}^{2}}. \label{DF}
\end{equation}
(where $F = A$ and $B$) denotes the standard deviations, $|\phi\rangle$ is a  state of the system under considerations, $\langle \psi|\phi\rangle$ is the inner product in the state space ${\cal H}$ of the system, ($|\psi\rangle, |\phi\rangle \in {\cal H}$), $\langle F\rangle_{\phi} = \langle \phi|F\,\phi\rangle \equiv \langle \phi|F|\phi\rangle$ is the average (expected) value of  the observable $F$ in the state $|\phi\rangle$, provided that $|\langle\phi|F|\phi \rangle |< \infty$.
 The observable $F$ is represented by hermitian operator $F$ acting in a Hilbert space ${\cal H}$ of states.
 Here  $[A,B] = AB - BA$  and it is assumed that  states $|\phi\rangle$ under consideration are  normalized: $\langle \phi|\phi\rangle \stackrel{\rm def}{=} \|\,|\phi\rangle \|^{2} =1$. (In Eq. (\ref{H1})  $F$ stands for position  and momentum operators $x$ and $p_{x}$).

	The inspiration for this paper comes from some recent publications in which the authors claim that the Heisenberg--Robertson (HR) uncertainty relation for two non--commuting observables,  $A$ and $B$,
has some imperfections
because it does not provide any information about the standard deviations when the system is in one of the eigenstates of one of these observables (see, eg. \cite{Mac,Hua,Bin,Xiao}).
	 It is claimed in these papers that much better uncertainty relations are the new ones of the type of  "sum uncertainty relations", because the above--mentioned "weak point" of the Heisenberg--Robertson relation does not appear  in these new ones. For this reason, in the following parts of the paper we will analyze in detail the properties of  uncertainty relations of this  and related types, and in particular we will examine their properties in the cases when the quantum system is in one of the eigenstates of the analyzed pair of non--commuting observables.
Also the properties of uncertainty relations of this type in other critical cases, i.e. in cases where the right hand side of the HR uncertainty relation  is equal to zero, will be analyzed in the further parts of this paper.

The paper is organized as follows: Sec. 2  contains, for the convenience of readers,  a  derivation of the Heisenberg--Robertson uncertainty relation and analysis of some its properties. In Sec. 3 readers will find selected inequalities for vectors from the state space ${\cal H}$, which are useful in studying the uncertainty principles. The derivation of the  uncertainty relation for the sum of variances and similar related uncertainty relations  is given in Sec. 4.
In Sec. 5, we analyze the critical  points of the uncertainty relations derived in Sec. 4. Discussion of the results presented in this paper and final remarks are given in Sec. 6.

\section{Heisenberg--Robertson uncertainty relation}
As we have already mentioned, the inequality (\ref{R1}) holds for each pair of non-commuting observables $A$ and $B$ represented by Hermitian operators acting in the state space ${\cal H}$ of the physical system under study,
such that $[A,B]$ exists and $|\phi\rangle \in {\cal D}(AB) \bigcap {\cal D}(BA) \subset {\cal H}$, (${\cal D}({\cal O})$ denotes the domain of an operator $\cal O$ or of a product of operators).
Note that, the relation (\ref{R1}) results from basic assumptions of the quantum theory and from the geometry of Hilbert space. Indeed,
taking into account that
in the general case for an observable $F$ the standard deviation is defined as follows
\begin{equation}
\Delta_{\phi} F = \|\delta_{\phi} F|\phi\rangle\| \geq 0, \label{dF}
\end{equation}
where  $\delta_{\phi} F = F - \langle F\rangle_{\phi}\,\mathbb{I}$,
and defining
\begin{equation}
\delta_{\phi} A = A - \langle A\rangle_{\phi}\,\mathbb{I},\;\;{\rm and,} \;\;
\delta_{\phi} B = B - \langle B\rangle_{\phi}\,\mathbb{I},
\end{equation}
 one finds that
\begin{equation}
[A,B] \equiv [\delta_{\phi} A, \delta_{\phi} B],
\end{equation}
for all $|\phi\rangle \in {\cal D}(AB) \bigcap {\cal D}(BA)$.

Taking all this into account, we get  that
\begin{eqnarray}
\left|\langle \phi| [A,B] |\phi \rangle \right|^{2} &\equiv & \left|\langle \phi| [\delta_{\phi} A,\delta_{\phi} B] |\phi \rangle \right|^{2} \label{[dA,dB]} \\
& \equiv & \left| \,\langle \phi| \delta_{\phi} A \;\delta_{\phi} B |\phi \rangle - \langle \phi| \delta_{\phi} B\; \delta_{\phi} A |\phi \rangle \,\right|^{2} \nonumber \\
& =  &  \left|\, \langle \phi| \delta_{\phi} A \;\delta_{\phi} B |\phi \rangle  \, - \, \left( \langle \phi| \delta_{\phi} A\; \delta_{\phi} B |\phi \rangle \right)^{\ast} \right|^{2} \nonumber \\
& =  & 4\,\left|\,{\rm Im.}\left[\langle \phi| \delta_{\phi} A \;\delta_{\phi} B |\phi \rangle \right]\, \right|^{2} \label{Im[]} \\
&\leq& 4 \,\left| \,\langle \phi| \delta_{\phi} A \;\delta_{\phi} B |\phi \rangle \,\right|^{2} \label{C00} \\
& \leq & 4 \,\left\| \delta_{\phi} A |\phi \rangle \right \|^{2}\,\cdot\,\left \|\delta_{\phi} B|\phi \rangle \right\|^{2}  \nonumber \\
& \equiv & 4 \;(\Delta_{\phi}A)^{2}\,\cdot \,(\Delta_{\phi}B)^{2}, \label{H-ku}
\end{eqnarray}
which reproduces inequality (\ref{R1}).  There is $ (\Delta_{\phi} A)^{2} \equiv \left\|\delta_{\phi} A|\phi\rangle \right\|^{2}$ and $ (\Delta_{\phi} B)^{2} \equiv \left\|\delta_{\phi} B|\phi\rangle \right\|^{2}$, (Relations (\ref{dF}) and (\ref{DF}) are equivalent). Here the Schwartz inequality was used:
\begin{equation}
\left| \,\langle \phi| \delta_{\phi} A \;\delta_{\phi} B |\phi \rangle \,\right|^{2} \leq  \,\left\| \delta_{\phi} A |\phi \rangle \right \|^{2}\,\cdot\,\left \|\delta_{\phi} B|\phi \rangle \right\|^{2}.
\label{Schw1}
\end{equation}

	There are physicists who consider the weak side of inequality (\ref{R1}) to be that its right--hand side is equal to zero when it is calculated for a vector $|\phi\rangle$ being  an eigenvector of one of the observables $A$ or $B$, which makes this inequality trivial, (see, e.g. \cite{Mac,Hua,Bin,Xiao}).
	Indeed,  if to
 assume that $[A,B]\neq 0$ and $|\phi \rangle = |\psi_{b}\rangle$ is a normalized eigenvector of  $B$ for the eigenvalue $b$, that is that
  \begin{equation}
  B|\psi_{b}\rangle = b |\psi_{b}\rangle. \label{B=b}
  \end{equation}
  then one  finds that $\langle \psi_{b}|AB|\psi_{b}\rangle = b \,\langle \psi_{b}|A|\psi_{b}\rangle$ and thus
  \begin{eqnarray}
  	\langle [A,B] \rangle_{\psi_{b}} &\equiv& \langle \psi_{b}|  [A,B] |\psi_{b} \rangle \nonumber \\
  	&=& \langle \psi_{b}|  AB |\psi_{b} \rangle - \langle \psi_{b}|  BA |\psi_{b} \rangle \equiv 0, \label{[A,B]=0}
  \end{eqnarray}
i.e., for  $|\phi\rangle = |\psi_{b}\rangle$, the right--hand side of (\ref{R1}) vanishes.
The  assumption (\ref{B=b}) immediately implies also that
  \begin{equation}
  \langle \psi_{b}|B|\psi_{b} \rangle = b,\; \;{\rm and,}\;\; \delta_{\psi_{b}} B|\psi_{b}\rangle \equiv 0. \label{B=b+1}
  \end{equation}
   As a result we have that
   \begin{equation}
 \Delta_{\psi_{b}}B = \left\| \delta_{\psi_{b}} B |\psi_{b} \rangle \right \| =0, \label{DBb=0}
 \end{equation}
which, together with (\ref{[A,B]=0}),  means that if $|\phi\rangle$ is an eigenstate of the observalbe $B$, $|\phi\rangle = |\psi_{b}\rangle$, then  we have no any information from (\ref{R1}) on $\Delta_{\psi_{b}}A$.

\section{Preliminaries: useful inequalities}

As a "cure" for the above described "weakness" of the uncertainty relation (\ref{R1}) a new uncertainty relation for the sum of the variances of two non-commuting observables, $A$ and $B$, has been proposed (see e.g. \cite{Mac,Pat} and also \cite{Hua,Hof} and other papers).  The starting points for the "sum uncertainty relations" are some inequalities taking place in the Hilbert state space.
The basis of various types of "sum uncertainty relations" are: the triangle inequality,
\begin{equation}
\left\|\;| \psi_{A}\rangle \right\| + \left\|\;| \psi_{B}\rangle \right\|\;\geq \; \left\|\;| \psi_{A}\rangle \; + \;| \psi_{B}\rangle \right\|,\label{t1}
\end{equation}
the triangle inequality of the second kind (see, e.g. \cite{Jam}),
\begin{equation}
\left\|\;| \psi_{A}\rangle \right\|^{2} + \left\|\;| \psi_{B}\rangle \right\|^{2}\;\geq \; \frac{1}{2}\left\|\;| \psi_{A}\rangle \; + \;| \psi_{B}\rangle \right\|^{2},\label{t2}
\end{equation}
which follows from the parallelogram law,
\begin{equation}
 2\left(  {\left\| \;|\psi_{A}  \rangle \right\|}^{2} +
{ \left\|\; |\psi_{B}  \rangle \right\|}^{2} \right)  =  { \left\|\; |\psi_{A}  \rangle +  |\psi_{B}  \rangle \right\| }^{2} + { \left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|}^{2}, \label{par}
\end{equation}
and other related inequalities. Let us note that from  (\ref{par}) there follows another inequality,
\begin{equation}
\left\|\;| \psi_{A}\rangle \right\|^{2} + \left\|\;| \psi_{B}\rangle \right\|^{2}\;\geq \; \frac{1}{2}\left\|\;| \psi_{A}\rangle \; - \;| \psi_{B}\rangle \right\|^{2}.\label{t2a}
\end{equation}
Yet another useful inequality
can be easily derived using the triangle inequality and the parallelogram law \cite{ku-rev1}.  Namely,
let's take the inequality (\ref{t1}) and square its both sides. We get that
\begin{equation}
\left\|\;| \psi_{A}\rangle \right\|^{2} + \left\|\;| \psi_{B}\rangle \right\|^{2} \,+\,2\,\left\|\;| \psi_{A}\rangle \right\|\left\|\;| \psi_{A}\rangle \right\|\;
\geq \; \left\|\;| \psi_{A}\rangle \; + \;| \psi_{B}\rangle \right\|^{2}.\label{t1-2}
\end{equation}
In the next step we find $ \left\|\;| \psi_{A}\rangle \; + \;| \psi_{B}\rangle \right\|^{2} $ from (\ref{par} ) and replace the right hand-side of (\ref{t1-2})  with the $ \left\|\;| \psi_{A}\rangle \; + \;| \psi_{B}\rangle \right\|^{2} $ calculated in this way,
which leads to the following inequality,
\begin{equation}
2\,\left\|\;| \psi_{A}\rangle \right\|\left\|\;| \psi_{A}\rangle \right\|
 \;\geq \; \left\|\;| \psi_{A}\rangle \right\|^{2}\, + \, \left\|\;| \psi_{B}\rangle \right\|^{2}\,
- \,  \left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|^{2}. \label{t1^2<}
\end{equation}
In  this way, by transforming (\ref{t1^2<}) suitably, we obtain the following  inequality,
\begin{equation}
\left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|^{2}\,+\,
2\,\left\|\;| \psi_{A}\rangle \right\|\left\|\;| \psi_{A}\rangle \right\|
 \;\geq \; \left\|\;| \psi_{A}\rangle \right\|^{2}\, + \, \left\|\;| \psi_{B}\rangle \right\|^{2}.  \label{new-ineq}
\end{equation}

Note that the reverse triangle inequality immediately follows from inequality (\ref{new-ineq}). Indeed, using (\ref{t1^2<}), (\ref{new-ineq}) one finds that
\begin{equation}
\left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|^{2}\;
 \;\geq \; \left\|\;| \psi_{A}\rangle \right\|^{2}\, + \, \left\|\;| \psi_{B}\rangle \right\|^{2}\,-\,
2\,\left\|\;| \psi_{A}\rangle \right\|\left\|\;| \psi_{A}\rangle \right\|,
  \label{rev1}
\end{equation}
that is that
\begin{equation}
\left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|^{2}\;
 \;\geq \;\left( \left\|\;| \psi_{A}\rangle \right\|\, - \, \left\|\;| \psi_{B}\rangle \right\|\right)^{2},
  \label{rev2}
\end{equation}
which is equivalent to the reverse triangle inequality:
\begin{equation}
\left\| \;|\psi_{A}  \rangle -  |\psi_{B}  \rangle \right\|\;
 \;\geq \;\left| \;{\left\|\;| \psi_{A}\rangle \right\|\, - \, \left\|\;| \psi_{B}\rangle \right\|}\;\right|.
  \label{rev3}
\end{equation}

Next, we can use the following property
\begin{equation}
\left(\left\|\,|\psi_{A}\rangle\right\| - \left\|\,|\psi_{B}\rangle \right\|\right)^{2} = \left\|\,|\psi_{A}\rangle\right\|^{2} + \left\|\,|\psi_{B}\rangle\right\|^{2} \,- \,2\left\|\,|\psi_{A}\rangle\right\|\, \left\|\,|\psi_{B}\rangle\right\|\,\geq\, 0. \label{n1}
\end{equation}
and obtain another inequality valid in the Hilbert space of states,
\begin{equation}
\left\|\,|\psi_{A}\rangle\right\|^{2} + \left\|\,|\psi_{B}\rangle\right\|^{2}\;\geq \;2\left\|\,|\psi_{A}\rangle\right\|\, \left\|\,|\psi_{B}\rangle\right\|.  \label{n2}
\end{equation}

Continuing our considerations, let us note that if we use the following identities,
\begin{eqnarray}
\left\|\,|\psi_{A}\rangle\, - \,|\psi_{B}\rangle\right\|^{2} &=&\left\|\,|\psi_{A}\rangle\right\|^{2}\,+\,\left\|\,|\psi_{B}\rangle\right\|^{2}\,
-\,2 \,\Re(\langle \psi_{A}|\psi_{B}\rangle), \label{id1}\\
\left\|\,|\psi_{A}\rangle\, + \,|\psi_{B}\rangle\right\|^{2} &=&\left\|\,|\psi_{A}\rangle\right\|^{2}\,+\,\left\|\,|\psi_{B}\rangle\right\|^{2}\,
+\,2 \,\Re(\langle \psi_{A}|\psi_{B}\rangle), \label{id2}
\end{eqnarray}
and then multiply them by themselves on their respective sides,
we obtain that,
\begin{eqnarray}
\left\|\,|\psi_{A}\rangle - |\psi_{B}\rangle\right\|^{2}\;\left\|\,|\psi_{A}\rangle + |\psi_{B}\rangle\right\|^{2} & = &
  \left(\left\|\,|\psi_{A}\rangle\right\|^{2}+\left\|\,|\psi_{B}\rangle\right\|^{2}\right)^{2}\nonumber  \\
  & &\;\;\;\;\;\;\;\;-4\;\left[\Re(\langle \psi_{A}|\psi_{B}\rangle)\right]^{2}.
\end{eqnarray}
Using this result we get finally (see, e.g., \cite{Bar}),
\begin{equation}
\left\|\,|\psi_{A}\rangle - |\psi_{B}\rangle\right\|\;\left\|\,|\psi_{A}\rangle + |\psi_{B}\rangle\right\|\ \leq
\left\|\,|\psi_{A}\rangle\right\|^{2}+\left\|\,|\psi_{B}\rangle\right\|^{2}. \label{Bar1}
\end{equation}

It is also worth mentioning here the generalization of the triangle inequality to $N$ vectors,
(see, e. g.,  \cite{Jam,Cer,Hsu}),
\begin{equation}
\sum_{j=1}^{N}\left\|\,|\psi_{j}\rangle \right\|\;\geq\; \left\|\sum_{j=1}^{N}|\psi_{j}\rangle \right\|, \label{t-n}
\end{equation}
which can be used to derive the "sum uncertainty relation" for $N$ non--commuting observables.

\section{Sum and other uncertainty relations}

Let us now derive the "sum uncertainty relation" and analyze some of its properties.
Inserting into (\ref{t1})
\begin{equation}
|\psi_{A}\rangle = \delta_{\phi} A|\phi\rangle,\;\; |\psi_{B}\rangle  = \delta_{\phi} B|\phi\rangle, \label{dA}
\end{equation}
 one gets
 \begin{equation}
 \left\|\;\delta_{\phi} A| \phi\rangle \right\| + \left\|\;\delta_{\phi} B| \phi\rangle \right\|\;\geq \; \left\|\delta_{\phi} A| \phi\rangle \; + \;\delta_{\phi} B| \phi\rangle \right\|. \label{P1}
 \end{equation}
 Next, taking into account  (\ref{dF}), we see that the inequality (\ref{P1}) can be written in terms of standard deviations as
\begin{equation}
\Delta_{\phi}A + \Delta_{\phi}B \geq \Delta_{\phi}(A + B). \label{Pat4}
\end{equation}
It is because simply,
\begin{equation}
\delta_{\phi} A| \phi\rangle \; \pm \;\delta_{\phi} B| \phi\rangle \equiv \delta_{\phi}( A \pm B)| \phi\rangle. \label{d(A+B)}
\end{equation}
The inequality (\ref{Pat4}) is the {\em sum uncertainty relation} derived in \cite{Pat} and discussed in the literature.

The generalization of the  triangle inequality
to $N$ vectors
allows generalizing the sum uncertainty relation (\ref{Pat4}) for two noncommuting observables to the case of $N$ non--commuting observables $\left\{ A_{j} \right\}_{j=1}^{N}$ with $ N \geq 2$ for standard deviations, $ (\Delta_{\phi}A_{1}), (\Delta_{\phi}A_{2}), \ldots , (\Delta_{\phi}A_{N})$,
\begin{equation}
\sum_{j=1}^{N}(\Delta_{\phi}A_{j}) \geq   \,\Delta_{\phi}(\sum_{j=1}^{N} A_{j}). \label{t-n1}
\end{equation}
This type of "sum uncertainty relation" and its various generalizations  have recently been intensively studied in many papers (see, e. g. \cite{Pat,Bin,Xiao,Son} and many others).


In addition to the "sum uncertainty relations" for standard deviations, the "stronger sum uncertainty relations" for variances have also been studied recently.
Such uncertainty relations can be obtained from  the triangle inequality of the second kind (\ref{t2}). Indeed,
starting from formula (\ref{t2}) and using (\ref{dA}) we get
\begin{equation}
\left\|\;\delta_{\phi} A| \phi\rangle \right\|^{2} + \left\|\;\delta_{\phi} B| \phi\rangle \right\|^{2}\;\geq \; \frac{1}{2}\left\|\delta_{\phi} A| \phi\rangle \; + \;\delta_{\phi} B| \phi\rangle \right\|^{2}.\label{m1}
\end{equation}
Next, taking into account   (\ref{d(A+B)}),
we can replace inequality (\ref{m1}) by the following one,
\begin{equation}
\left\|\;\delta_{\phi} A| \phi\rangle \right\|^{2} + \left\|\;\delta_{\phi} B| \phi\rangle \right\|^{2}\;\geq \; \frac{1}{2}\left\|\;|\delta_{\phi}( A + B)| \phi\rangle \right\|^{2}.\label{m2}
\end{equation}
Keeping in mind (\ref{dF})  one concludes that the inequality (\ref{m2}) means that
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \geq	 \frac{1}{2} \left(\Delta_{\phi}(A+B)\right)^{2}, \label{M12}
\end{equation}
and this inequality is called "the stronger sum uncertainty relations" and  is
the one of sum of variances uncertainty relations derived in  \cite{Mac}.

Note that there is also a related relation resulting from (\ref{t2a}). Repeating all the steps leading to the inequality (\ref{M12}) we obtain,
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \geq	 \frac{1}{2} \left(\Delta_{\phi}(A-B)\right)^{2}. \label{M12aa}
\end{equation}
This relation was studied, eg. in \cite{strong-2}. In the same way, using (\ref{new-ineq}), we find that
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2}  \leq \left( \Delta_{\phi}(A-B) \right)^{2} + 2 \Delta_{\phi} A\; \Delta_{\phi} B, \label{rev1a}
\end{equation}
which is a new form of the reverse uncertainty relation derived and discussed in \cite{Deb,Xiao2,Xiao3} and in other papers.

The method used to derive the above uncertainty relations and applied to other inequalities discussed in Sec. 3 allows us to derive other, similar relations. Thus, from the reverse triangle inequality (\ref{rev3}) we obtain,
\begin{equation}
\Delta_{\phi}(A - B)
 \;\geq \;\left|\Delta_{\phi}A\, - \, \Delta_{\phi}B\right|.  \label{n4}
\end{equation}
One can conclude from (\ref{n2}) that (see also \cite{Mac,strong-2}),
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \;\geq\;2\,\Delta_{\phi}A\;\Delta_{\phi}B, \label{n5}
\end{equation}
whereas
from (\ref{Bar1}) that
\begin{equation}
\Delta_{\phi}(A-B)\;\Delta_{\phi}(A+B) \leq \left(\Delta_{\phi}A\right)^{2} +  \left(\Delta_{\phi}B\right)^{2}. \label{Bar2}
\end{equation}
This last sum uncertainty relation was derived in \cite{Bar}.

To conclude this Section, it is worth noting that the inequality (\ref{n5}) can be interpreted in two ways: once as  another lower bound on the sum of variances, and secondly as an upper bound on the product of standard deviations, which product appears in the Heisenberg--Robertson  uncertainty principle (\ref{R1}).
This second possibility means that we have found a  reverse uncertainty relation for the Heisenberg--Robertson uncertainty relation.

\section{Critical points  of sum and other uncertainty relations}

Now let's check what happens with the inequality (\ref{Pat4}) when $|\phi\rangle$ is an eigenstate, e.g. of the operator $B$, i.e. when $|\phi\rangle = |\psi_{b}\rangle$.
 In such a case, relations (\ref{B=b}) and (\ref{B=b+1}) are satisfied,  which means that property  (\ref{DBb=0}) takes place,  and by virtue of (\ref{d(A+B)}) we obtain
\begin{equation}
\delta_{\psi_{b}}( A \pm B) | \psi_{b} \rangle  \equiv \delta_{\psi_{b}} A | \psi_{b} \rangle  \pm \delta_{\psi_{b}} B| \psi_{b}\rangle \equiv
 \delta_{\psi_{b}} A| \psi_{b}\rangle. \label{dA+dB=dA}
\end{equation}
The elementary conclusion following from this last relation is that the right hand--side of the inequality (\ref{Pat4}) calculated for $|\phi\rangle = |\psi_{b}\rangle$ looks as follows
\begin{equation}
\Delta_{\psi_{b}}(A + B) \equiv  \Delta_{\psi_{b}} A, \label{DA+DB=DA}
\end{equation}
and consequently the inequality (\ref{Pat4}) takes the form,
\begin{equation}
 \Delta_{\psi_{b}} A \geq   \Delta_{\psi_{b}} A. \label{Pat4=0}
\end{equation}
Thus, if the quantum system under study is in a state that is an eigenstate of one of the observables, e.g. $B$, then, similarly to the HR uncertainty relation, the sum uncertainty relation (\ref{Pat4}) does not provide any information about a lower bound on the standard deviation $ \Delta_{\psi_{b}} A$ of the observable $A$.


We will now focus on the analysis of the properties of the inequality (\ref{M12}),
or more precisely, on the study of the equivalent inequality (\ref{m2}) preceding it.
Let us now repeat our calculations performed for the Heisenberg-- Robertson uncertainty principle and for the sum uncertainty relation (\ref{Pat4})
in the case where $|\phi\rangle$ is an eigenvector of the operator, e. g. $B$.
Again, using results
(\ref{B=b}), (\ref{B=b+1}),  (\ref{DBb=0}) and (\ref{d(A+B)}) we can conclude that the result (\ref{DA+DB=DA}) takes also place in the considered case.
This in turn means that in this case the inequality (\ref{M12})
  takes the following form (see also, e.g., \cite{ku-edu})
\begin{equation}
	(\Delta_{\psi_{b}} A)^{2}  \geq	 \frac{1}{2} \left( \Delta_{\psi_{b}} A \right)^{2}. \label{ku1}
\end{equation}
Remember, it was assumed that
$ [A,B] \neq 0 $, therefore
$A$ and  $B$ can not have common eigenvectors and thus there must be
$\Delta_{\psi_{b}}(A) >0$.
As a result we have  two absolutely trivial solutions of the inequality  (\ref{ku1}).
The first one: $ 1 > 1/2$, and the second one:
\begin{equation}
	(\Delta_{\psi_{b}}A)^{2}  \geq	0. \label{DAb>0}
\end{equation}
Again, the last result is the same as the one that, which follows from the Heisenberg--Robertson uncertainty principle.


We see that, in the case when the system is in state $|\phi\rangle$ which is an eigenstate of, for example, operator $B$, $|\phi\rangle = |\psi_{b}\rangle$, in contrast to the HR uncertainty relation (\ref{R1}), (see (\ref{[A,B]=0})),  the right-hand sides of the sum uncertainty relations (\ref{Pat4=0}) and (\ref{ku1}) are non--zero. However, this does not mean that these relations are non-trivial. Simply, in this case these sum uncertainty relations are trivially fulfilled for any $\Delta_{\psi_{b}}A \geq 0$.


Now let's check what happens with inequalities  (\ref{rev1a}) and (\ref{n4})  when the state $|\phi\rangle$ is an eigenstate of operator $B$, i.e. when  $|\phi\rangle = |\psi_{b}\rangle$. Proceeding in this case in the same way as in the case of (\ref{Pat4}) and (\ref{M12}) we get
\begin{equation}
(\Delta_{\psi_{b}}A)^{2} \geq (\Delta_{\psi_{b}}A)^{2},
\end{equation}
for inequality (\ref{rev1a}), and
\begin{equation}
(\Delta_{\psi_{b}}A) \geq \left|\Delta_{\psi_{b}}A\right|
\end{equation}
for inequality (\ref{n4}). (Here  results (\ref{dA+dB=dA}), (\ref{DA+DB=DA}) and (\ref{DBb=0}) were used). Similar results can be obtained for the inequality (\ref{Bar2}). The above procedure applied to the inequality (\ref{Bar2}) leads to the following result,
\begin{equation}
(\Delta_{\psi_{b}}A)^{2} \leq (\Delta_{\psi_{b}}A)^{2}. \label{Bar3}
\end{equation}
As can be seen from this, in the case when the state $|\phi\rangle$ is an eigenstate of one of the two observables, $A$ and $B$,  e.g. $B$, from inequalities (\ref{rev1a}), (\ref{n4}) and (\ref{Bar2}) we also do not obtain any significant bounds on the standard deviation, e. g. $\Delta_{\psi_{b}}A$, of the observable $A$. Similar results take place  when $|\phi\rangle$  is an eigenvector of $A$.


It turns out that in the state space one can find vectors that are not eigenvectors of any pair of non--commuting observables $ A$ and $B$, and yet the uncertainty relations analyzed above
do not provide any nontrivial lower or upper bounds on the sums of standard deviations and variances calculated for these vectors, (see, e.g. \cite{ku-min,ku1}).
This property is true when vectors $|\psi_{A}\rangle = \delta_{\phi}A|\phi\rangle$ and $|\psi_{B}\rangle = \delta_{\phi}B|\phi\rangle $  are non--zero and orthogonal to each other: $\langle \psi_{A}|\psi_{B}\rangle \equiv
\langle \phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle =0$ and $|\phi\rangle$ is not an eigenvector of either $A$ or $B$.
Then $\Delta_{\phi} A > 0$ and $\Delta_{\phi}B > 0$.
The product
\begin{equation}
{\cal C}_{\phi}(A,B) \stackrel{\rm def}{=} \langle \phi|\delta_{\phi}A\;\delta_{\phi}B|\phi\rangle \equiv \langle AB \rangle_{\phi}\, -\, \langle A\rangle_{\phi}\,\langle B\rangle_{\phi}, \label{C1}
\end{equation}
is a quantum version of the correlation function (see, e. g.,  \cite{Poz,Khr} and references in \cite{ku-min}). The property  ${\cal C}_{\phi}(A,B) = 0$  means that  observables $A$ and $B$ are uncorrelated in the state $|\phi\rangle$.
So, if ${\cal C}_{\phi}(A,B) =0$ then
$ |\psi_{A}\rangle \perp |\psi_{B}\rangle$
 and in such a case  we have
 \begin{equation}
 \left\| \psi_{A} \rangle \pm |\psi_{B}\rangle \right\|^{2} \equiv \left\| \,|\psi_{A} \rangle \right\|^{2} + \left\|\,|\psi_{B}\rangle \right\|^{2},\;\; {\rm if}\;\; |\psi_{A}\rangle \perp |\psi_{B}\rangle.
 \label{perp}
\end{equation}
If we now assume that $ |\psi_{A}\rangle \perp |\psi_{B}\rangle$ and square both sides of the triangle inequality (\ref{t1}) and apply the result (\ref{perp}), then instead of the inequality (\ref{t1}) we simply get that
\begin{equation}
2\,\left\| \,|\psi_{A} \rangle \right\|^{2} \; \left\|\,|\psi_{B}\rangle \right\|^{2} \geq 0. \label{perp2}
\end{equation}
Then, remembering the definitions (\ref{dA}), from the last inequality we conclude that the sum uncertainty relation (\ref{Pat4}) in the considered case simplifies to
\begin{equation}
\Delta_{\phi}A\,\cdot\,\Delta_{\phi}B \,\geq\,0.  \label{t-hr}
\end{equation}
Applying the result (\ref{perp}) to the triangle inequality of the second kind (\ref{t2}) and proceeding as in the previous case, we find that the stronger sum uncertainty relations (\ref{M12}) takes the following form
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2}\,\geq \,0, \label{perp+t2}
\end{equation}
in the considered case.
The common feature of inequalities (\ref{t-hr}), (\ref{perp+t2}) is that they do not provide any significant information about the lower bound on the sum of standard deviations or on the sum of variations. Simply, the information resulting from these results is the same as that known from the definition (\ref{DF}) of the standard deviation $\Delta_{\phi}F$ of the observable $F$.
Similar conclusions hold for observables $A$ and $ B$ uncorrelated in state $|\phi\rangle$ in the case of the remaining uncertainty relations discussed in the previous section.

Note that in the case where,  observables $A$ and $B$ are uncorrelated in state $|\phi\rangle$, i.e. when $|\psi_{A}\rangle = \delta_{\phi}A|\phi\rangle \;\perp\;|\psi_{B}\rangle = \delta_{\phi}B|\phi\rangle $, then equations (\ref{[dA,dB]}) and  (\ref{C00}) imply that $|\langle \phi|[A,B]|\phi\rangle| = 0$.
This property means that that if ${\cal C}_{\phi}(A,B) = 0$, then the right hand--side of the standard HR uncertainty relation (\ref{R1}) is equal  to zero despite the fact that $\Delta_{\phi}A  > 0$  and $\Delta_{\phi}B >0$.
To perform a more in--depth analysis of this property, let us note that the following relation follows from (\ref{[dA,dB]}),  (\ref{Im[]}) and (\ref{C1}),
\begin{equation}
\langle \phi| [A,B] |\phi \rangle \equiv 2 \Im\,[{\cal C}_{\phi}(A,B)]. \label{[A,B]=ImC}
\end{equation}
So, in fact, as can be seen from (\ref{Im[]}) and (\ref{[A,B]=ImC}), it is enough that $\Im\,[ C_{\phi}(A,B)=0$ so that $|\langle \phi|[A,B]|\phi\rangle| = 0$ in the inequality (\ref{R1}).
It should be noted here that in such cases, i.e. when $\Im\,[ {\cal C}_{\phi}(A,B)]= 0$, the stronger sum uncertainty relation (\ref{M12}) has some advantages over the standard Heisenberg--Robertson uncertainty relation (\ref{R1}). Namely, using equations (\ref{id1}), (\ref{id2}), (\ref{m1}) and (\ref{m2}), (or simply using (\ref{n2}) and  (\ref{n5})), one can  show that the inequality (\ref{M12}) is equivalent to the following one (see \cite{ku-edu}),
\begin{equation}
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \geq	2|{\cal C}_{\phi}^{clas}(A,B)|, \label{M12+new}
\end{equation}
where ${\cal C}_{\phi}^{clas}(A,B) \stackrel{\rm def}{=} \Re\,[ C_{\phi}(A,B)]$ is the classical part of $C_{\phi}(A,B)$ and corresponds to the classical correlation function.
Thus, in all those cases when $\Im\,[{\cal C}_{\phi}(A,B)] = 0$, which results in that the right side of the HR inequality (\ref{R1}) is equal to zero, the right side of the stronger uncertainty relation (\ref{M12}) in the form (\ref{M12+new}) can be different from zero.  So, even in such cases, contrary to the standard HR uncertainty relation,   there can be a lower, non--zero, bound for the sum of variances $(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2}$.
On the other hand, the inequality (\ref{M12+new})  is simply an upper bound on the modulus of the real part of the quantum version of correlation function ${\cal C}_{\phi}(A,B)$, i.e., on modulus of ${\cal C}_{\phi}^{clas}(A,B)$.

\section{Final remarks}

The Heisenberg--Robertson  uncertainty relation (\ref{R1}) as well as sum uncertainty relations (\ref{Pat4}), (\ref{M12}) and others derived above were be proven using mathematically exact inequalities resulting from the properties of Hilbert spaces.
There are many rigorous inequalities in Hilbert space \cite{Cer} which can be used to find various
relations  of the kind discussed above.
Many of them provide us with information about the lower bound on the product or sum of standard deviations, or variances and these lower bounds are studied in many papers.
At this point it is worth mentioning that the properties of Hilbert spaces provide us with the tools to find not only these lower bounds, but also other relations and bounds for the sums of standard deviations or variances.
For example, by inserting $|\psi_{A}\rangle$  and $|\psi_{B}\rangle$ in the form (\ref{dA}) into the parallelogram law (\ref{par}),
 we find that there must be:
\begin{equation}
2\left( (\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \right) = \left(\Delta_{\phi}(A + B)\right)^{2} + \left((\Delta_{\phi}(A - B)\right )^{2}. \label{DA+DB-5}
\end{equation}
Another example:
In \cite{Deb} it was shown  using the Dunkle-Williams inequality \cite{Cer,Dun} that there is an upper bound on the sum of variances. A similar bound is expressed by the inequality (\ref{rev1a}) derived above.
The upper bound (\ref{rev1a}) is the consequence of the triangle inequality and of the parallelogram law, which means that it is  rigorous one. So, the quantum physics, strictly speaking,  the geometry of Hilbert space of states of a quantum system forces the sum of variances to have both the lower (\ref{M12}) and the upper (\ref{rev1a}) bounds. In the literature, inequalities of the type of (\ref{rev1a}) are called the reverse uncertainty relations (see, e.g. \cite{Deb,Xiao2,Xiao3,Xiao4}). It seems worth noting that the above derivation of the reverse uncertainty relation based on our inequality (\ref{new-ineq}) is much simpler than those found in the literature \cite{Deb,Xiao2,Xiao3,Xiao4} and based on the  Dunkl--Williams inequality \cite{Cer,Dun}.

As  mentioned earlier,  there is another upper bound:  The inequality (\ref{n5}) can be considered as an upper bound for the product of standard deviations $\Delta_{\phi}A\;\Delta_{\phi}B$, that is as an upper bound for the Heisenberg--Robertson uncertainty relation (see also \cite{Mac}),
\begin{equation}
\frac{1}{2}\left( (\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \right) \;\geq \;\Delta_{\phi}A\;\Delta_{\phi}B\; \geq\;\frac{1}{2} \left|\langle [A,B] \rangle_{\phi} \right|. \label{n6}
\end{equation}
In other words,
inequalities (\ref{n5}) and (\ref{n6}) show that the Heisenberg--Robertson uncertainty relation also has its reverse uncertainty relation.

Similarly, an analogous chain of inequalities can be easily found for the sum of variances. Namely, combining inequalities (\ref{rev1a}) and (\ref{M12}), we find
\begin{equation}
\left( \Delta_{\phi}(A-B) \right)^{2} + 2 \Delta_{\phi} A\; \Delta_{\phi} B\,\geq \,
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \, \geq  \,  \frac{1}{2}\,  [\Delta_{\phi}(A +B)]^{2}. \label{nv}
\end{equation}
On the other hand, if to use relations (\ref{rev1a}) and (\ref{n5}) one obtains
\begin{equation}
\left( \Delta_{\phi}(A-B) \right)^{2} + 2 \Delta_{\phi} A\; \Delta_{\phi} B\,\geq \,
(\Delta_{\phi}A)^{2} + (\Delta_{\phi}B)^{2} \, \geq  \,   2 \Delta_{\phi}A\;\Delta_{\phi}B. \label{nv1}
\end{equation}

Note that
some of the results presented in Sec. 4, 5 and 6 point to a rather unexpected aspect of the quantum world. Namely, the conclusions resulting from the uncertainty relation HR (\ref{R1}), or from the relation for the sum of variances (see  (\ref{M12}), (\ref{Bar2})), say that there are lower bounds on the product of standard deviations or on the sum of variances. They mean that the spread of the expected values of non--commuting observables $A$ and $B$ in state $|\phi\rangle$ cannot be too small. On the other hand, the conclusions resulting from the reverse uncertainty relations, saying that there are upper bounds on the product of standard deviations  (\ref{n6}), or on the sum of variances (\ref{rev1a}), can be considered as a surprise. In other words, in the quantum world the spread of the values of the non--commuting observables $A$ and $B$ measured in state $|\phi\rangle$ can not be
as small as one wishes, but also it can not
too large, which is a rather unexpected conclusion resulting from (\ref{n6}) and (\ref{rev1a}). (Similar conclusions results from the reverse uncertainty relation, which was  derived in \cite{Deb} using the  Dunkl--Williams inequality \cite{Cer,Dun}).

As mentioned, using various  inequalities valid in the Hilbert state space, one can find various relations between standard deviations or variances of non--commuting observables. These can be lower bounds on products of variances or standard deviations, or lower or upper bounds on their sums. They can also be relations similar to Eq. (\ref{DA+DB-5}). It seems, that relations of the type discussed above may be useful, e.g., in quantum metrology. In general, their usefulness, as well as the usefulness of the relations discussed in Sec. 4, depends on the goal we want to achieve, or the phenomena we want to study, describe  and understand. For this reason, all relations of the mentioned types require further research.

\section*{Acknowledgments}
This work was supported by
the program of the Polish Ministry
of Science and Higher Education under the name "Regional
Initiative of Excellence", Project No. RID/SP/0050/2024/1.

\end{document}